# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

### v1

@@ Abstract
In recent years, graph neural networks (GNNs) have become a popular tool to improve the accuracy and performance of recommender systems. Modern recommender systems are not only designed to serve the end users, but also to benefit other participants, such as items and items providers. These participants may have different or conflicting goals and interests, which raise the need for fairness and popularity bias considerations. GNN-based recommendation methods also face the challenges of unfairness and popularity bias and their normalization and aggregation processes suffer from these challenges. In this paper, we propose a fair GNN-based recommender system, called HetroFair, to improve items’ side fairness. HetroFair uses two separate components to generate fairness-aware embeddings: i) fairness-aware attention which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees, and ii) heterophily feature weighting to assign distinct weights to different features during the aggregation process. In order to evaluate the effectiveness of HetroFair, we conduct extensive experiments over six real-world datasets. Our experimental results reveal that HetroFair not only alleviates the unfairness and popularity bias on the items’ side, but also achieves superior accuracy on the users’ side. Our implementation is publicly available at https://github.com/NematGH/HetroFair
@@ 1 Introduction
Graph-based recommendation has become more popular in recent years, due to the existence of a bipartite graph in the nature of recommender systems [1, 2, 3, 4, 5, 6]. In this bipartite graph, nodes are divided into two sets users and items, and edges are ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [7, 8, 9]. With the widespread use of graph neural networks in recommender systems, the competition to build systems with higher accuracy entered a new phase [2, 10, 11, 12]. Early methods attempted to improve the accuracy of recommender systems, to better satisfy end users. However, users are not the only stakeholders and accuracy is not the sole objective. Recent researches demonstrate that recommender systems are multi-stakeholder systems in which in addition to end users, other participants (especially items and items providers) should also be considered, during the recommendation process [13, 14, 15, 16, 17]. On the other hand, the power law characteristic of the recommendation graphs shows that a large number of items possess a low degree, while only a few of them have a large degree. Figure 1 illustrates the degree distribution of two real-world recommendation datasets. This feature causes a phenomenon, known in the literature as popularity bias, where popular items are recommended more frequent than the other items. As a result, items with few interactions (long-tail items) are ignored, even if they are of high quality and relevance [18, 19]. Similar investigations have been done for GNN-based recommendation, revealing that long-tail items are often sacrificed during message normalization and aggregation of GNNs [20, 21]. Various methods have been proposed in recent years for long-tail recommendation, which aim to recommend items that belong to the long tail of the distribution of ratings or popularity and at the same time, try to achieve high accuracy on the users’ side. The authors of [22] introduce regularization techniques to overcome the popularity bias. Some other works concentrate on creating multiple graphs for different objectives and simultaneously optimize those graphs in an integrated manner [23, 24]. A few other works directly target the normalization and aggregation processes in the recommendation systems that utilize graph neural networks [20, 21]. However, these GNN-based recommenders either consider static weights for message normalization, or ignore different aspects of items/users. In this paper, we propose the HetroFair model, to improve items’ side fairness in GNN-based recommendation systems. To do so, first we directly address the cause of popularity bias and reduce the effect of nodes’ degrees on enlarging the values of their embeddings, via involving dot products during the normalization process. Second, with respect to the heterophily property in graphs, we distinguish between different aspects of an item, by considering an specific weight for each feature during the aggregation process. In the heterophily property, nodes of the graph that are not similar to each other, tend to be connected to each other. To the best of our knowledge, our model is the first approach that addresses fairness in GNN-based recommendation using the heterophily property. Our key contributions are summarized as follows:
@@ 2.2 Fairness-aware recommendation
The primary objective of popularity debias is to unearth qualified but less interacted items, long-tail ones, which are items that have lower training samples than short-head items. V2HT [23] constructs a heterogeneous item-item graph with four different types of edges to connect long-tails to popular items. SGL [34] brings self-supervised learning to the recommendation setting and creates an auxiliary task for the long-tail recommendation. r-AdjNorm [20] changes the power of the symmetric square root normalization term in graph neural networks to control the normalization process during neighborhood aggregation, aiming for improved results, especially for low-degree items. APDA [21] assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. Our proposed method takes a different approach to improve items’ side fairness and on the one hand, considers different weights for features during the aggregation process. On the other hand, it directly involves dot products in the normalization process.
@@ 3.2 Graph neural networks
Here, $D$ is the diagonal degree matrix that has the same dimension as the adjacency matrix $A$ of the graph, $D_{ii}=\sum_{j}A_{ij}$, $\tilde{D}=D+I$, $\tilde{A}=A+I$, $I\in R^{\left(|U|+|I|\right)\times\left(|U|+|I|\right)}$ is the identity matrix, $\sigma\left(\cdot\right)$ is an activation function, $H^{\left(k\right)}$ is the matrix of node embeddings (representations) in the layer $k$ and $M^{\left(k\right)}$ denotes a layer-specific trainable weight matrix.
@@ 3.3 Light Graph Convolutional
Here, $h_{u}^{\left(k\right)}$ and $h_{i}^{\left(k\right)}$ denote the embeddings of user $u$ and item $i$ at layer $k$. By $N\left(x\right)$ we denote the set of nodes which have an edge to node $x$, and by $d_{x}$ we denote the degree of node $x$. The term $\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}$ is called the symmetric square root normalization term. In the matrix form, $H^{\left(0\right)}=[h_{1}^{\left(0\right)},h_{2}^{\left(0\right)},\ldots,h_{|V|+|E|}^{\left(0\right)}]$ denotes the initial embeddings layer. The embeddings in the next layers are formulated as follows:
@@ 4.1 Symmetric square root normalization effect
$y_{ui}=h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i^{\prime}}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2}.$ (10)
@@ 4.2 Fairness-aware attention
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. By incorporating the sigmoid of dot products alongside the previous normalization term, messages from long-tail items (such as $i$) receive a higher importance coefficient than short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients. This not only prevents the over-enlargement of representations but also ensures that the effects of large representations are gradually transmitted through the computational graph neighborhood.
@@ 4.3 Heterophily feature weighting
Statistics #Nodes #Edges #Classes Homophily
@@ 4.4 Model training
where $O_{u}^{+}$ is user intracted items during training and $T=\left\{\left(u,i,j\right)\mid i\in O_{u}^{+}\wedge j\in I\setminus O_{u}^{+}\right\}$ indicates the training set. The hyperparameter $\beta$ controls regularization to prevent overfitting, $||\Theta||$ is the model’s parameters that must be learned during the optimization process, $\tilde{y}_{ui}$ and $\tilde{y}_{uj}$ are user interest scores for items $i$ and $j$ respectively, obtained through the dot product of the embeddings of the user and item.
@@ 4.5 The algorithm
In this section, we present the process of fair embedding generation, along with the algorithmic steps that our recommendation algorithm follows during a batch. Algorithm 1 describes one batch training process, in our model. First, in Lines 3-6, we initialize model parameters using the Xavier initialization method [38]. The embedding matrix, updated during the training process, is also initialized. In Line 9, we call the FairEmbeddingGeneration function, presented in Algorithm 2 to calculate users’ and items’ embeddings. Subsequently, for every user-item pair in the current batch $O_{batch}$, we choose a negative sample uniformly at random with replacement, as explained in [37]. In Lines 12-15, we compute the interest score of the user for positive and negative items, calculate the loss, and finally backpropagate the error to update model parameters. Algorithm 2 outlines the fair embedding generation process. For every node in the graph and for each layer, we act as the follow. First in Line 7, we compute the fairness-aware attention as the dot products of the embeddings of users and items, for all items in the users’ training profile. Then, we transform these attention weights into a new vector space to consider distinct attention for each feature, as shown in Line 8. In Line 10, by taking the inverse of the weight in the message passing process, we enhance the likelihood of suggesting long-tail items to users while simultaneously diminishing the dominance of popular items. Finally, in Line 12, we combine the embeddings from different layers to obtain the final representations of items/users.
@@
It is imperative for recommendation systems to deliver not only satisfactory performance but also operate efficiency. Striking a balance between these two criteria is often viewed as a trade-off. In this context, we delve into the analysis of the time complexity of our proposed model. If we neglect Lines 6-9 of Algorithm 2, our model becomes the same as the LightGCN model, which has the time complexity of $O\left(|E|Kd\right)$ in the graph convolution part, where $K$ and $d$ are constants [39]. Thanks to parameter sharing, the primary operation in Equation 13, which appears in Lines 7-8, is the multiplication of a $1\times 1$ matrix with a $1\times d$ matrix, which can be done in $O\left(d\right)$ time. Hence, the additional time complexity that our proposed method imposes on the basic LightGCN model is $O\left(d\right)$. According to this analysis, the time complexity of our proposed graph convolution process is $O\left(|E|\left(Kd+d\right)\right)=O\left(|E|Kd\right)$, which is the same as the LightGCN model.
@@ 5.1.2 Baselines
Datasets #User #Item #Intraction Density
@@ 5.1.3 Evaluation metrics
Here, $REL_{N}$ is a sorted list of the $N$ most relevant items in ascending order, $r(i)$ is the relevance of the $i^{th}$ ranked item, which is either 00 or $1$ respectively for irrelevant and relevant items. MRR is concerned with the position of the first relevant item:
@@ 5.1.4 Hyper-parameters setting
For all the methods, we use the same early stopping mechanism, stopping training after $15$ epochs if NGCF does not show improvement in these $15$ epochs. The embeddings dimension is set to $64$ for LightGCN and AdjNorm, $256$ for the APDA model, and $128$ for our model. The only exception is the Movies dataset where due to hardware limitations, we set the embeddings dimension to 64. The batch size for LightGCN and AdjNorm is set to $1024$, as proposed in their papers, while for APDA we set it to $2048$. The learning rate for all the baselines is set to $0.001$. For all the methods, we set both the regularization hyper-parameter $\beta$ and the weight decay parameter to $0.0001$. The learning rate for our model is set to $0.0005$. We set the number of propagation layers to $3$ in LightGCN and r-AdjNorm, and to $4$ in APDA and HetroFair.
@@ 5.2 Performance comparison
Table 3 shows the results of HetroFair in comparison to the other methods. The findings indicate that our method outperforms the other methods, not only in the fairness metrics but also in the accuracy metrics. These results show that our method is better fitted with the data and takes into account their distinct features.
@@
Dataset Metric LightGCN r-AdjNorm APDA HetroFair w/o Hetro
Epinions NDCG 0.0789 0.0787 0.0859 0.0895 0.0889
MRR 0.1361 0.1347 0.1471 0.1525 0.1522
MAP 0.0329 0.0328 0.0366 0.0379 0.0377
PRU 0.5233 0.5310 0.5188 0.4706 0.5054
PRI 0.5129 0.5404 0.4983 0.4384 0.4573
Electronics NDCG 0.0463 0.0440 0.0517 0.0525 0.0488
MRR 0.0645 0.0608 0.0708 0.0733 0.0676
MAP 0.0221 0.0207 0.0250 0.0256 0.0232
PRU 0.5216 0.5469 0.5548 0.4600 0.5020
PRI 0.5417 0.5834 0.5692 0.4462 0.4805
CDs NDCG 0.1231 0.1219 0.1379 0.1449 0.1346
MRR 0.1715 0.1609 0.1940 0.2017 0.1882
MAP 0.0611 0.0588 0.0710 0.0747 0.0680
PRU 0.3304 0.3366 0.3265 0.2972 0.3266
PRI 0.2824 0.3297 0.2763 0.2264 0.2495
Health NDCG 0.1259 0.1283 0.1314 0.1334 0.1351
MRR 0.1986 0.2040 0.2073 0.2112 0.2093
MAP 0.0595 0.0614 0.0637 0.0656 0.0669
PRU 0.5162 0.5330 0.5205 0.4922 0.4830
PRI 0.2164 0.2088 0.2266 0.1821 0.1270
Beauty NDCG 0.2113 0.2184 0.2274 0.2308 0.2302
MRR 0.2525 0.2616 0.2790 0.2824 0.2798
MAP 0.1200 0.1267 0.1354 0.1364 0.1389
PRU 0.4428 0.4221 0.4418 0.4139 0.3117
PRI 0.2164 0.2088 0.2266 0.1821 0.1270
Movies NDCG 0.0672 0.0681 0.0715 0.0777 0.0741
MRR 0.0966 0.0982 0.1011 0.1093 0.1048
MAP 0.0304 0.0308 0.0331 0.0365 0.0340
PRU 0.3098 0.3404 0.3384 0.3045 0.3251
PRI 0.3951 0.4545 0.3862 0.3355 0.3590

### v2

@@ Abstract
In recent years, graph neural networks (GNNs) have become a popular tool to improve the accuracy and performance of recommender systems. Modern recommender systems are not only designed to serve end users, but also to benefit other participants, such as items and items providers. These participants may have different or conflicting goals and interests, which raise the need for fairness and popularity bias considerations. GNN-based recommendation methods also face the challenges of unfairness and popularity bias and their normalization and aggregation processes suffer from these challenges. In this paper, we propose a fair GNN-based recommender system, called HetroFair, to improve items’ side fairness. HetroFair uses two separate components to generate fairness-aware embeddings: i) fairness-aware attention which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees, and ii) heterophily feature weighting to assign distinct weights to different features during the aggregation process. In order to evaluate the effectiveness of HetroFair, we conduct extensive experiments over six real-world datasets. Our experimental results reveal that HetroFair not only alleviates the unfairness and popularity bias on items’ side, but also achieves superior accuracy on users’ side. Our implementation is publicly available at https://github.com/NematGH/HetroFair.
@@ 1 Introduction
Graph-based recommendation has become more popular in recent years, due to the existence of a bipartite graph in the nature of recommender systems [5, 14, 25, 29, 30, 32]. In this bipartite graph, nodes are divided into two sets of users and items, and edges are ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [43, 10, 6, 45, 19]. With the widespread use of graph neural networks in recommender systems, the competition to build systems with higher accuracy entered a new phase [14, 9, 18, 37]. Early methods attempted to improve the accuracy of recommender systems, to better satisfy end users. However, users are not the only stakeholders and accuracy is not the sole objective. Recent researches demonstrate that recommender systems are multi-stakeholder systems in which in addition to end users, other participants (especially items and items providers) should also be considered, during the recommendation process [1, 2, 4, 21, 23]. On the other hand, the power law characteristic of the recommendation graphs shows that a large number of items possess a low degree, while only a few of them have a large degree. Figure 1 illustrates the degree distribution of two real-world recommendation datasets. This feature causes a phenomenon, known in the literature as popularity bias, where popular items are recommended more frequent than the other items. As a result, items with few interactions (long-tail items) are ignored, even if they are of high quality and relevance [3, 20]. Similar investigations have been done for GNN-based recommendation, revealing that long-tail items are often sacrificed during message normalization and aggregation of GNNs [39, 42].
Various methods have been proposed in recent years for long-tail recommendation, which aim to recommend items that belong to the long tail of the distribution of ratings or popularity and at the same time, try to achieve high accuracy on the users’ side. The authors of [44] introduce regularization techniques to overcome the popularity bias. Some other works concentrate on creating multiple graphs for different objectives and simultaneously optimize those graphs in an integrated manner [17, 28]. A few other works directly target the normalization and aggregation processes in the recommendation systems that utilize graph neural networks [39, 42]. However, these GNN-based recommenders either consider static weights for message normalization, or ignore different aspects of items/users.
In this paper, we propose the HetroFair model, to improve items’ side fairness in GNN-based recommendation systems. To do so, first we directly address the cause of popularity bias and reduce the effect of nodes’ degrees on enlarging the values of their embeddings, via involving dot products during the normalization process. Second, with respect to the heterophily property in graphs, we distinguish between different aspects of an item, by considering an specific weight for each feature during the aggregation process. In the heterophily property, nodes of the graph that are not similar to each other, tend to be connected to each other. To the best of our knowledge, our model is the first approach that addresses fairness in GNN-based recommendation using the heterophily property. Our key contributions are summarized as follows:
@@ 2.2 Fairness-aware recommendation
The primary objective of popularity debias is to unearth qualified but less interacted items, long-tail ones, which are items that have lower training samples than short-head items. V2HT [17] constructs a heterogeneous item-item graph with four different types of edges to connect long-tails to popular items. SGL [31] brings self-supervised learning to the recommendation setting and creates an auxiliary task for the long-tail recommendation. Zhao et al. [40] disentangle popularity bias into two benign and harmful categories and argue that item quality can also affect popularity bias. They propose a time-aware disentangled framework to identify harmful bias. r-AdjNorm [39] changes the power of the symmetric square root normalization term in graph neural networks to control the normalization process during neighborhood aggregation, aiming for improved results, especially for low-degree items. APDA [42] assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. Chen et al. [7] theoretically analyze the effect of graph convolution on popularity bias amplification and show that stacked multiple graph convolution layers accelerate users’ movement towards popular items in the representation space. Our proposed method takes a different approach to improve items’ side fairness and on the one hand, considers different weights for features during the aggregation process. On the other hand, it directly involves dot products in the normalization process.
@@ 3.2 Graph neural networks
Here, $D$ is the diagonal degree matrix that has the same dimension as the adjacency matrix $A$ of the graph, $D_{ii}=\sum_{j}A_{ij}$, $\tilde{D}=D+I$, $\tilde{A}=A+I$, $I\in R^{\left(|U|+|I|\right)\times\left(|U|+|I|\right)}$ is the identity matrix, $\sigma\left(\cdot\right)$ is an activation function, $H^{\left(k\right)}$ is the matrix of node embeddings (representations) in the layer $k$ and $M^{\left(k\right)}$ denotes a layer-specific trainable weight matrix. To compute the embedding of each node in the graph, a graph neural network constructs a graph, which is a rooted tree and is usually called node’s computational graph [33].
@@ 3.3 Light Graph Convolutional
Here, $h_{u}^{\left(k\right)}$ and $h_{i}^{\left(k\right)}$ denote the embeddings of user $u$ and item $i$ at layer $k$. By $N\left(x\right)$ we denote the set of nodes which have an edge to node $x$, and by $d_{x}$ we denote the degree of node $x$. The term $\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}$ is called the symmetric square root normalization term. In the matrix form, $H^{\left(0\right)}=[h_{1}^{\left(0\right)},h_{2}^{\left(0\right)},\ldots,h_{|V|}^{\left(0\right)}]$ denotes the initial embeddings layer. The embeddings in the next layers are formulated as follows:
@@ 4.1 Symmetric square root normalization effect
$y_{ui^{\prime}}=h_{u}^{\left(k\right)}\cdot h_{i^{\prime}}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i^{\prime}}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2}.$ (10)
@@ 4.2 Fairness-aware attention
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. Therefore, the result of the sigmoid function for long-tail items will be close to zero, and for short-head items, it will be close to one. By incorporating the sigmoid of dot products along with the above-mentioned normalization term, long-tail items achieve a smaller denominator than short-head items. This leads to messages from long-tail items (such as $i$) receiving a higher importance coefficient than those from short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients. This not only prevents the over-enlargement of representations, but also ensures that the effects of large representations are gradually transmitted through nodes’ computational graphs.
@@ 4.3 Heterophily feature weighting
Dataset #nodes #edges #classes Homophily
@@ 4.4 Model training
where $T=\left\{\left(u,i,j\right)\mid i\in O_{u}^{+}\wedge j\in I\setminus O_{u}^{+}\right\}$ indicates the training set and $O_{u}^{+}$ is user interacted items during training. The hyperparameter $\beta$ controls regularization to prevent overfitting, $\Theta$ is the model’s parameters that must be learned during the optimization process, $\tilde{y}_{ui}$ and $\tilde{y}_{uj}$ are user interest scores for items $i$ and $j$ respectively, obtained through the dot product of the embeddings of the user and item.
@@ 4.5 The algorithm
In this section, we present the process of fair embedding generation, along with the algorithmic steps that our recommendation algorithm follows during a batch. Algorithm 1 describes one batch training process, in our model. First, in Lines 3-6, we initialize model parameters using the Xavier initialization method [11]. The embedding matrix, updated during the training process, is also initialized. In Line 9, we call the FairEmbeddingGeneration function, presented in Algorithm 2 to calculate users’ and items’ embeddings. Subsequently, for every user-item pair in the current batch $O_{batch}$, we choose a negative sample uniformly at random with replacement, as explained in [24]. In Lines 12-15, we compute the interest score of the user for positive and negative items, calculate the loss, and finally backpropagate the error to update model parameters.
Algorithm 2 outlines the fair embedding generation process. For every node in the graph and for each layer, we act as the follow. First in Line 7, we compute the fairness-aware attention as the dot products of the embeddings of users and items, for all items in the users’ training profile. Then, we transform these attention weights into a new vector space to consider distinct attention for each feature, as shown in Line 8. In Line 10, by taking the inverse of the weight in the message passing process, we enhance the likelihood of suggesting long-tail items to users while simultaneously diminishing the dominance of popular items. Finally, in Line 12, we combine the embeddings from different layers to obtain the final representations of items/users.
@@
It is imperative for recommendation systems to deliver not only satisfactory performance but also operate efficiency. Striking a balance between these two criteria is often viewed as a trade-off. In this context, we delve into the analysis of the time complexity of our proposed model. If we neglect Lines 6-9 of Algorithm 2, our model becomes the same as the LightGCN model, which has the time complexity of $O\left(|E|Kd\right)$ in the graph convolution part, where $K$ and $d$ can be seen as constants [36]. Thanks to parameter sharing, the primary operation in Equation 13, which appears in Lines 7-8, is the multiplication of a $1\times 1$ matrix with a $1\times d$ matrix, which can be done in $O\left(d\right)$ time. Hence, the additional time complexity that our proposed method imposes on the basic LightGCN model is $O\left(d\right)$. According to this analysis, the time complexity of our proposed graph convolution process is $O\left(|E|\left(Kd+d\right)\right)=O\left(|E|Kd\right)$, which is the same as the LightGCN model.
@@ 5.1.1 Datasets
Dataset #users #items #interactions Density
@@
5.1.2 Baselines
  • Reg [44]: This is a regularization-based method that considers the correlation between items’ popularity and their predicted score by a model. Following the original paper, we tune the hyperparameter $\gamma$ carefully to achieve comparable results in accuracy and fairness.
  • PC [44]: This is a pos-processing (ReRanking) approach that manipulates predicted scores to calculate new scores as a function of item popularity. We tune the parameter $\alpha$ and $\beta$ as described in the original paper.
@@ 5.1.3 Evaluation metrics
Here, $REL_{N}$ is a sorted list of the $N$ most relevant items in ascending order, $r(i)$ is the relevance of the $i^{th}$ ranked item, which is either 00 or $1$ respectively for irrelevant and relevant items.
MRR is concerned with the position of the first relevant item:
@@ 5.1.4 Hyper-parameters setting
For all methods, we use the same early stopping mechanism, stopping training after $15$ epochs if NGCF does not show improvement in these $15$ epochs. The embeddings dimension is set to $64$ for LightGCN and AdjNorm, $256$ for the APDA model, and $128$ for our model. The only exception is the Movies dataset where due to hardware limitations, we set the embeddings dimension to 64. The batch size for LightGCN and AdjNorm is set to $1024$, as proposed in their papers, while for APDA we set it to $2048$. The learning rate for all the baselines is set to $0.001$. For all the methods, we set both the regularization hyper-parameter $\beta$ and the weight decay parameter to $0.0001$. The learning rate for our model is set to $0.0005$. We set the number of propagation layers to $3$ in LightGCN and r-AdjNorm, and to $4$ in APDA and HetroFair.
@@ 5.2 Performance comparison
Table 3 shows the results of HetroFair in comparison to the other methods. The best performance is marked in bold, and the second-best performance is underlined. The findings indicate that our method outperforms the other methods, not only in the fairness metrics but also in the accuracy metrics. These results show that our method is better fitted with the data and takes into account their distinct features. In the case of baselines, it is observable that although the Reg method often achieves commendable fairness metrics compared to other baselines, its accuracy is significantly lower than that of other methods. This suggests that regularization-based methods may reduce the performance of recommendations for end-users.
@@
Dataset Metric LightGCN LightGCN+Reg LightGCN+PC r-AdjNorm APDA Our model w/o Hetro
Epinions NDCG 0.0789 0.0746 0.0778 0.0798 0.0859 0.0895 0.0889
MRR 0.1361 0.1288 0.1336 0.1369 0.1471 0.1525 0.1522
MAP 0.0329 0.0307 0.0324 0.0339 0.0366 0.0379 0.0377
PRU 0.5233 0.5114 0.5168 0.5077 0.5188 0.4706 0.5054
PRI 0.5129 0.5065 0.5071 0.5366 0.4983 0.4384 0.4573
Electronics NDCG 0.0463 0.0407 0.0422 0.0461 0.0517 0.0525 0.0488
MRR 0.0645 0.0552 0.0610 0.0640 0.0708 0.0733 0.0676
MAP 0.0221 0.0188 0.0203 0.0219 0.0250 0.0256 0.0232
PRU 0.5216 0.4779 0.4919 0.5300 0.5548 0.4600 0.5020
PRI 0.5417 0.5077 0.5325 0.5661 0.5692 0.4462 0.4805
CDs NDCG 0.1231 0.1230 0.1245 0.1178 0.1379 0.1449 0.1346
MRR 0.1715 0.1723 0.1760 0.1638 0.1940 0.2017 0.1882
MAP 0.0611 0.0611 0.0617 0.0563 0.0710 0.0747 0.0680
PRU 0.3304 0.3287 0.3270 0.3262 0.3265 0.2972 0.3266
PRI 0.2824 0.3069 0.2974 0.3193 0.2763 0.2264 0.2495
Health NDCG 0.1259 0.1202 0.1224 0.1271 0.1314 0.1334 0.1351
MRR 0.1986 0.1902 0.196 0.2028 0.2073 0.2112 0.2093
MAP 0.0595 0.0570 0.0574 0.0610 0.0637 0.0656 0.0669
PRU 0.5162 0.4980 0.4893 0.5270 0.5205 0.4922 0.4830
PRI 0.5303 0.5116 0.5285 0.5240 0.5555 0.5028 0.4804
Beauty NDCG 0.2113 0.2161 0.2116 0.2136 0.2274 0.2308 0.2302
MRR 0.2525 0.2603 0.2540 0.2520 0.2790 0.2824 0.2798
MAP 0.1200 0.1240 0.1203 0.1235 0.1354 0.1364 0.1389
PRU 0.4428 0.4407 0.04385 0.4188 0.4418 0.4139 0.3117
PRI 0.2164 0.2071 0.2151 0.2089 0.2266 0.1821 0.1270
Movies NDCG 0.0672 0.0673 0.0690 0.0676 0.0715 0.0777 0.0741
MRR 0.0966 0.0957 0.0971 0.0955 0.1011 0.1093 0.1048
MAP 0.0304 0.0298 0.0308 0.0299 0.0331 0.0365 0.0340
PRU 0.3098 0.3051 0.3082 0.3245 0.3384 0.3045 0.3251
PRI 0.3951 0.3714 0.3829 0.4436 0.3862 0.3355 0.3590